OpenStudy (anonymous):

Help me with the rest of completing this square, please.. y = 2x^2 +10x -12 y+12= 2x^2+10x y+12= 2(x^2 +5x 5/2 = 2.5 &; 2.5^2= 6.25 y+12+12.5= 2(x^2 +5x +6.25) I'm stuck here.

OpenStudy (anonymous):

Im not sure i follow your work, want me to show you my steps and then compare?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So you're solving for x?

OpenStudy (anonymous):

No, That's why I deleted lol. I realized it's y=

OpenStudy (anonymous):

So you just want it in vertex form?

OpenStudy (anonymous):

general form.

OpenStudy (anonymous):

It's already in general form though...above you say complete the square y=a(x-h)^2+k that's vertex form

OpenStudy (anonymous):

Than in vertex form, I'm just going by what my paper is telling me. My next step should be factoring the trinomial (x^2 +5x +6.25) and then bringing the numbers on the left side of the equation back to the right so it is y=(the trinomial factored) -24.5. Does that make any sense, or.. no lol

OpenStudy (anonymous):

*equal sign

OpenStudy (anonymous):

Hmm, here ill show you the steps and then we can get an idea 2x^2+10x-12 -pull out the 2 2(x^2+5x-6) complete the square inside parentheses 2(x^2+5x+25/4-25/4-6) factor first trinomial in parentheses and combine like terms 2((x+5/2)^2-49/4) redistribute the 2 y=2(x+5/2)^2-49/2

OpenStudy (anonymous):

It's similar to the steps I have to follow bc I also get 5/2 & 5/2^2.. hints.. the 2.5 and the 6.25. Basically the only thing I need to complete my problem following these confusing behind steps that I have is (x^2 +5x +6.25) factored.

OpenStudy (anonymous):

You got 6.25 from dividing 5 by 2 and then squaring it once you do that and add the number in there it will as x plus or minus the number you had before squaring it (x+5/2)^2 or (x+2.5)^2

OpenStudy (anonymous):

I was just about to type that... y + 24.5 = 2(x + 2.5)^2 y = 2(x + 2.5)^2 - 24.5

OpenStudy (anonymous):

Thank you.